A horse is harnessed to a sled having a mass of 241 kg, including supplies. The horse must exert a force exceeding 1260 N at an angle of 30.3° (above the horizontal) in order to get the sled moving. Treat the sled as a point particle.




(a) Calculate the normal force (in N) on the sled when the magnitude of the applied force is 1260 N. (Enter the magnitude.) __N


(b) Find the coefficient of static friction between the sled and the ground beneath it.


(c) Find the static friction force (in N) when the horse is exerting a force of 6.30  102 N on the sled at the same angle. (Enter the magnitude.) __N



1
Expert's answer
2021-09-18T15:00:16-0400


(a)

"N +Fsinθ =mg \\

N= mg -Fsinθ \\

N = 241 \times 9.81 -1260 \times sin(30.3°) \\

N= 2364.21 -635.69 \\

N= 1728.52\;N"

(b) For static

"\sum fx = 0 \\

Fcosθ -f =0 \\

Fcosθ = \mu N \\

\mu = \frac{Fcosθ}{N} \\

\mu = \frac{1260 \times cos(30.3°)}{N} \\

\mu = \frac{1260 \times 0.86339 }{1728.52} \\

\mu = 0.629"

(c) If F = 630 N

"N = mg -Fsinθ \\

N = 241 \times 9.81 -630 \times sinθ \\

N = 2364.21 -317.84 \\

N = 2046.37 \;N \\

f_R = \mu N \\

f_R= 0.629 \times 2046.37 \\

f_R = 1287.16 \;N"